So I have an expression like this:
35: 3☆+1❐+0△
59: 1☆+2❐+1△
50: 2☆+2❐+0△
? : 2☆+1❐+1△
and I need to find the variable:?
I have tried many ways like:
35: 3☆+1❐+0△
59: 1☆+2❐+1△
50: 2☆+2❐+0△
? : 2☆+1❐+1△
70: 6☆+2❐+0△
354: 6☆+12❐+6△
150: 6☆+6❐+0△
3?: 6☆+3❐+3△
35: 3☆+1❐+0△
59: 1☆+1❐+½△
25: 1☆+1❐+0△
? : 2☆+1❐+1△

Please help find variable:?

Answer by Boreal(15235)

35: 3☆+1❐+0△
59: 1☆+2❐+1△
50: 2☆+2❐+0△
I am assuming that the number on the left equals the variables on the right.
I like dealing with x, y, and z
3x+y=35
2x+2y=50
-6x-2y=-70, multiplying the first by -2
-4x=-20
x=5
y=20
z=14 using these in the middle equation
2x+y+z=44, where x=star, y=square and z=triangle